2. Phosphorus (V) oxide P4O10 is used to prepare phosphoric acid.

a) Write the reaction equation when solid phosphorus (V) oxide is prepared by burning solid phosphorus P4.

b) How much phosphorus (V) oxide can be produced when 1.33 g of phosphorus and 5.07 g of oxygen are available?
1
Expert's answer
2020-12-28T04:41:45-0500

Solution:

a) The reaction equation when solid phosphorus (V) oxide is prepared by burning solid phosphorus P4:

P4(s) + 5O2(g) → P4O10(s)

According to the equation: n(P4) = n(O2)/5 = n(P4O10)


b):

1) Moles of P4 = Mass of P4 / Molar mass of P4

The molar mass of P4 is 123.895 g mol-1.

Hence,

Moles of P4 = 1.33 g / 123.895 g mol-1 = 0.0107 mol

n(P4) = 0.0107 moles

2) Moles of O2 = Mass of O2 / Molar mass of O2

The molar mass of O2 is 15.999 g mol-1.

Hence,

Moles of O2 = 5.07 g / 15.999 g mol-1 = 0.3169 mol

n(O2) = 0.3169 moles


According to stoichiometry:

1 mol of P4 reacts with 5 mol of O2

Thus 0.0107 moles of phosphorus (P4) reacts with:

(0.0107 mol P4 × 5 mol O2) / 1 mol P4 = 0.0535 mol O2.

However, initially there is 0.3169 moles of O2 (according to the task).

Thus P4 acts as limiting reagent and O2 is excess reagent.


Therefore,

n(P4O10) = n(P4) = 0.0107 mol (according to the equation).


The molar mass of P4O10 is 283.886 g mol-1.

Mass of P4O10 = Moles of P4O10 × Molar mass of P4O10

Mass of P4O10 = 0.0107 mol × 283.886 g mol-1 = 3.0376 g = 3.04 g

Mass of P4O10 is 3.04 grams.


Answer:

a) The balanced chemical equation: P4(s) + 5O2(g) → P4O10(s)

b) 3.04 grams of P4O10 (or 0.0107 mol P4O10) can be produced.